Hanna:

Suppose a researcher is studying a population that has a mean of μ=38.6 and a standard deviation of σ=1.1. The researcher studies a simple random sample of size n = 200. According to the central limit theorem, what is the approximate standard deviation of the sampling distribution of the mean? Round the answer to the nearest thousandth if needed. 0.006 0.078 0.193 2.73

mhchen:

So you use this formula: \[\frac{\sigma}{\sqrt{n}} = \frac{1.1}{\sqrt{200}}\]

Hanna:

Thank you I'm sure its 0.078